# Frontiers of Generative AI for Network Optimization: Theories, Limits, and Visions

### v1

@@ II-B Generative AI
GenAI aims to learn the probabilistic distribution of data, i.e., to model the joint probability $P(x,y)$ to be able to synthesize new samples in the original distribution. In contrast, discriminative AI only focuses on conditional probability $P(y|x)$, which is used to complete discriminative tasks such as classification and regression. Generative models are originally designed for multi-modal generation tasks such as images, texts, and audio. And generative models are capable of approximating conditional probability distributions when guided by conditions or prompts.
@@ II-B1 GDMs
GDMs have emerged as one of the most prominent generative frameworks in recent years. Initially designed for image synthesis, GDMs have since been extended to a wide array of domains, including the generation of solutions for complex optimization problems. The widely acknowledged starting point of this research line is the DDPM [22], which progressively adds Gaussian noise to data samples and trains a U-Net architecture [57] to reverse this process and reconstruct clean outputs. Building on this foundation, Classifier-Guidance [58] and classifier-free guidance (CFG) [59] introduced two of the most influential mechanisms for conditional generation. A substantial body of work has since been dedicated to improving the efficiency of GDMs, including denoising diffusion implicit model (DDIM) [60] and DPM-Solver [61], along with subsequent developments such as [62, 63, 64]. Breakthroughs in high-resolution image synthesis were achieved with Stable Diffusion [65], which introduced the latent diffusion paradigm. Meanwhile, ControlNet [66] significantly enhanced the controllability and quality of conditional image generation. Beyond continuous data domains, GDMs have also been extended to support discrete data generation [67]. Additionally, the backbone neural architectures have evolved to include advanced alternatives such as graph transformer (GAT) [68] and diffusion transformer (DiT) [69], reflecting growing architectural diversity and task specialization.
@@ II-B2 LPTMs
LLMs are the most important part of LPTMs. LLMs represent another major category of generative models, represented by GPT [23], LLaMA [70], Gemini [71], DeepSeek [72], etc. They typically adopt an autoregressive approach to model the conditional probability $P(x_{t}|x_{<t})$ at each token position in a sequence, where $x_{t}$ represents the $t$th token in the output sequence. This allows them to generate a wide range of outputs, including natural language text, programming code, structured instructions, and more.
Some of the latest foundation models have evolved from traditional LLMs into what are now called large reasoning models, such as DeepSeek-R1 [73]. These models are typically fine-tuned on top of LLMs using reinforcement learning to promote step-by-step reasoning or tree-search-like problem-solving behaviors, often with a focus on enabling reflection and self-correction. DeepSeek-R1, for example, is built using group relative policy optimization (GRPO) training guided by feedback from DeepSeek-V3 [72]’s judgment feedback. Closely related are mathematical LLMs, such as rStar-Math [74], DeepSeek-Prover-V2 [75], MINIMO [76], and LLMOPT [77], which specialize in mathematical reasoning tasks. These models often adapt general reasoning methods to math problems, commonly incorporating inference-time scaling (a strategy akin to extending thinking time in humans) and leveraging tree search to improve logical depth.
In parallel, some models aim to fundamentally rethink the architecture and training strategies of LLMs, particularly through diffusion-based language models and world models. LLaDA [78] and MMaDA [79] use diffusion processes instead of autoregressive generation, enabling controllable output lengths, denoising-style masked generation, and multi-modal input-output support. World models, on the other hand, aim to learn latent representations of physical environments and predict object dynamics within them. Notably, the latest V-JEPA 2 [80] is pre-trained on large-scale unlabeled video data and further fine-tuned on a small set of robotic control videos to enable V-JEPA 2 AC to support zero-shot robotic control.
These recent advancements in reasoning models, diffusion language models, and world models hold promising implications for network optimization tasks. However, how to effectively adapt and tailor them to networking scenarios remains an open challenge. In this stage, LPTMs have been increasingly applied in diverse scenarios such as mathematical modeling [81], optimization problem analysis [82], and even approximate problem-solving [83].
@@ II-B3 Others
Beyond images and text, GenAI has continued to advance across various multi-modal domains, including image-text contrastive learning (CLIP [84]), visual segmentation (SAM [85]), tabular data (CARTE [86]), graph-structured data (GraphEdit [87]), and multi-modal composition (CoDi [88]).
@@ II-C Neural Combinatorial Optimization
Combinatorial optimization refers to the class of problems where the goal is to find the optimal configuration or sequence from a discrete and often combinatorially large solution space, under given constraints. Classical examples include the traveling salesman problem (TSP), maximum independent set (MIS), and vehicle routing, all of which are NP-hard and difficult to scale with traditional solvers. Combinatorial optimization can be viewed as a class of problems encompassed by network optimization.
Neural combinatorial optimization aims to tackle these problems using neural networks to learn heuristics or direct solution mappings. GDMs has recently emerged as a powerful and popular approach for solving challenging combinatorial optimization problems [89, 90, 91, 92, 93]. These studies primarily focus on large-scale instances of classical combinatorial optimization problems, often involving hundreds of thousands of nodes, and aim to develop pre-trained models on canonical problem types that can generalize effectively to problem variants. Unlike reinforcement learning-based solvers [94], GDMs can directly model the solution distribution and generate near-optimal candidates in a single forward pass, making them well-suited for one-shot optimization scenarios common in network settings. This growing body of work has offered valuable inspiration for how GenAI, especially GDMs, can be applied to generate high-quality solutions in network optimization contexts.
Figure 4: Comparison between GDM and LLM in terms of generation mechanism. The sources of randomness in GDMs and LLMs differ: in GDMs, randomness arises from the initial noise sample and the stochastic nature of the denoising trajectory, whereas in LLMs, it stems from sampling predicted tokens. Moreover, LLMs are autoregressive, as each predicted token is fed back into the model for subsequent predictions, while most GDMs are non-autoregressive.
@@ II-D Deep Reinforcement Learning
DRL combines the perception capabilities of deep learning with the decision-making capabilities of reinforcement learning. It involves learning optimal policies through interactions between an agent and its environment to maximize cumulative rewards. By utilizing deep neural networks as function approximators, DRL is capable of addressing complex decision-making problems [95, 96].
GDM is one of the earliest GenAI applied in DRL, showcasing significant potential in robot control applications [97, 98, 99, 100, 101, 102, 103, 104, 105, 106]. Its unique capabilities in modeling complex data distributions have since enabled gradual expansion to broader DRL tasks. Recent advancements highlight GDM’s effectiveness in trajectory generation, where it provides smooth and diverse motion paths for agents [99]. In multi-task learning scenarios, GDM facilitates knowledge transfer across different tasks, improving overall learning efficiency [106]. Moreover, through data augmentation, GDM enhances the quality and diversity of training datasets, which is crucial for improving the generalization ability of DRL agents [100]. The application of GDM in DRL is being widely migrated to the MDP framework of network optimization modeling [40]. Ongoing innovations in the AI community are driving performance improvements in MDP-based network optimization within the networking domain.
@@ III GDMs in Network Optimization
Most applications of GDMs in network optimization are inspired by advances in the AI community. These applications primarily fall into two categories: using GDMs as solution generators for combinatorial optimization problems [89, 90, 91, 92, 93], and using GDMs as policy networks in DRL for robotic control [97, 98, 99, 100, 101, 102, 103, 104, 105, 106], as shown in Table IV.
When used as solution generators for combinatorial optimization, GDMs are typically discrete diffusion models targeting large-scale problems (often involving over 100,000 nodes). In the AI domain, GDMs have achieved near-optimal performance for classical combinatorial optimization tasks like TSP and MIS. Recent efforts have focused on accelerating solution generation [90] and enabling cross-problem generalization by leveraging shared knowledge across problem instances [93]. In the DRL setting, GDMs are commonly applied to generate action-state trajectories for robotic control. In robotic control, GDMs are mostly used in offline reinforcement learning, where they learn to predict the target trajectory given the history trajectory from the empirical data on interaction with the environment. Most existing approaches only execute the first action of the generated trajectory, which leads to inefficient exploration and significant time overhead. Current research is therefore exploring ways to refine generated trajectories with planning to improve action generation efficiency [102, 105]. General-purpose DRL using GDMs has also been explored as a promising research direction [100, 101, 106].
@@ III-A GDMs as Solution Generators
Network optimization problems are inherently more diverse and complex than classical combinatorial optimization. They encompass not only discrete-variable problems but also non-convex optimization, mixed-integer programming, and others. Compared to traditional combinatorial tasks, network optimization problems often involve more intricate objective functions and constraints, as they must account for various factors such as connectivity, latency, and total available resources.
@@
GDMOPT [40] is one of the earliest works to apply GDMs as solution generators in the network domain, conducting a case study on a classical convex optimization problem. DiffSG [107] demonstrates the effectiveness and cross-domain generalization of GDMs on a variety of more complex network optimization tasks. [108] further provides a theoretical explanation of why GenAI can outperform discriminative models in solution generation, offering insights into GDMs’ underlying mechanisms and analyzing the impact of key hyperparameters in network optimization settings, with DiffSG as a baseline. DFSS [109] aims to address the security of channel state information (CSI)-based sensing in integrated sensing and communication (ISAC) scenarios. It tackles unauthorized sensing or spoofing attacks by adversaries trying to exploit CSI signals. DFSS first uses a discrete graph diffusion model to generate connection topologies based on the deployment of ISAC devices, minimizing operational cost while preserving sensing performance. Then, it applies a continuous diffusion model to generate safeguarding CSI values, embedding them into pilot signals to significantly reduce the sensing accuracy of unauthorized devices.
@@
GDSG [110] considers the common challenge in network optimization where optimal solutions are difficult to obtain. It extends GDM training to support robust learning from sub-optimal solution samples only. Trained solely on efficiently constructed sub-optimal datasets, GDSG successfully optimizes multi-server multi-user computation offloading problems in multi-access edge computing (MEC) scenarios, involving discrete diffusion for offloading decisions and continuous diffusion for resource allocation. DEITSP [111] focuses on reducing the computational complexity of autoregressive GDM solvers. DEITSP employes a one-step diffusion mechanism that enables inference from any timestep toward the target solution. It also introduces a novel modification to the DDIM [60] process: each solution candidate undergoes a one-step prediction followed by re-noising and reintegration into the denoising process to enhance exploration. Additionally, it streamlines the GAT backbone to reduce inference time while maintaining solution quality, achieving faster results compared to DIFUSCO [89] and T2T [91]. GDPlan [112] addresses network planning by incorporating an external energy score-based model to guide graph diffusion for internet protocol (IP) link capacity configuration. High-quality solutions are further refined through external solvers in GDPlan.
@@
TABLE IV: Representative works of GDM as a solution generator and DRL policy in the AI community.
{tblr}
colspec=X[0.14,l] X[0.22,c] X[0.15,c] X[0.63,c], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, cell21 = r=5, cell71 = r=10, Type & Algorithm Neural Backbone Contributions
@@
Energy-Guided Sampling [93] GCN Proposed a general energy-guided sampling framework that enables a pre-trained GDM for the TSP to generate solutions across different scales and TSP variants.
@@
IDQL [98] MLP Demonstrated that the commonly used simple Gaussian policy struggles to perform well in learning Q-function solely from offline datasets under implicit Q-learning (IQL) settings. At the same time, GDMs can effectively take over this role with their powerful multi-modal distribution learning capability.
@@
MTDIFF [100] Transformer Leveraged GDM for both trajectory planning and data synthesis in multi-task DRL.
@@
EDP [102] MLP Designed an action approximation technique that enables direct reconstruction of actions during training without executing the full sampling chain combined with the DPM-Solver, significantly improving training efficiency.
SRDP [103] MLP Considered out-of-distribution (OOD) generalization issues in offline reinforcement learning caused by mismatch between real-world states and the training distribution by using GDM to learn generalizable state reconstruction features.
@@ III-B GDMs as DRL Policy Networks
Many network optimization problems can be modeled as MDP problems, but they differ from traditional robotic control tasks in two significant ways. First, the reward functions (or objective functions) and constraints in network optimization are typically far more complex than those in robotic control. While robot control tasks often have clear terminal states, making it easier to learn critic modules, network optimization problems often involve non-differentiable objectives and constraints, which cannot be directly optimized and must be approximated via external relaxation. These surrogate modules are sensitive to the objective formulation and hyperparameter settings. Second, in terms of real-time performance, dynamic network optimization typically demands decision generation within milliseconds to avoid system failures. In contrast, the actions generated by GDMs in robotic control usually reside in higher-level planning layers, with the real-time control handled at lower levels, thereby relaxing the real-time constraints on GDMs.
@@
DLRS [122] addresses the use of vector databases for request caching in edge-deployed LLMs. Observing that existing methods overlook previous search results, [122] proposes a new MDP-based quality of service (QoS) optimization framework called DLRS, improving both similar and novel request response frequencies in a cloud-edge cooperative setting. In [123], a GDM serves as a contract generator in a downlink semantic communication system, jointly optimizing rate-splitting multiple access (RSMA)-based resource allocation and contract design, achieving higher rewards than traditional DRL and supporting downstream semantic segmentation tasks. AC-GDM [124] targets secure beamforming in a multi-user IoT system with imperfect channels, using GDM-based DRL to jointly optimize precoding at the base station and phase shifts at an intelligent reflecting surface (IRS), outperforming traditional methods in simulation. GAI-Asynchronous [125] considers a low-altitude multi-task over-the-air edge inference system, modeling the joint optimization of task batching and beamforming, and uses GDMs as DRL policies to perform batch-wise beamforming decisions.
In summary, GDMs demonstrate superior performance as DRL policy networks in MDP-formulated network optimization problems, thanks to their powerful capability in modeling complex, multi-modal action distributions. However, as in the broader AI community, the high latency of GDMs due to their iterative denoising process remains a limitation. Furthermore, the complexity of network optimization objectives and constraints presents substantial challenges for developing general-purpose GDM-based DRL approaches and achieving cross-problem generalization in the network domain.

### v2

@@ II-B Generative AI
GenAI aims to learn the probabilistic distribution of data, i.e., to model the joint probability $P(\textbf{x},\textbf{y})$ to be able to synthesize new samples in the original distribution. In contrast, discriminative AI only focuses on conditional probability $P(\textbf{y}|\textbf{x})$, which is used to complete discriminative tasks such as classification and regression. Generative models are originally designed for multimodal generation tasks, such as images, text, and audio. And generative models can approximate conditional probability distributions when guided by conditions or prompts.
@@ II-B1 GDMs
GDMs have emerged as one of the most prominent generative frameworks in recent years. Initially designed for image synthesis, GDMs have since been extended to a wide array of domains, including the generation of solutions for complex optimization problems. The widely acknowledged starting point of this research line is the DDPM [22], which progressively adds Gaussian noise to data samples and trains a U-Net model [57] to reverse this process and reconstruct clean outputs. Building on this foundation, Classifier-Guidance [58] and classifier-free guidance (CFG) [59] introduced two of the most influential mechanisms for conditional generation. A substantial body of work has since been dedicated to improving the efficiency of GDMs, including denoising diffusion implicit model (DDIM) [60] and DPM-Solver [61], along with subsequent developments such as [62, 63, 64]. Breakthroughs in high-resolution image synthesis were achieved with Stable Diffusion [65], which introduced the latent diffusion paradigm. Meanwhile, ControlNet [66] significantly enhanced the controllability and quality of conditional image generation. Beyond continuous data domains, GDMs have also been extended to support discrete data generation [67]. Additionally, the backbone neural architectures have evolved to include advanced alternatives such as graph transformer (GAT) [68] and diffusion transformer (DiT) [69], reflecting growing architectural diversity and task specialization.
@@ II-B2 LPTMs
LLMs are the most important part of LPTMs, represented by GPT [23], LLaMA [70], Gemini [71], DeepSeek [72], etc. They typically adopt an autoregressive approach to model the conditional probability $P(\mathbf{x_{t}}|\mathbf{x_{<t}})$ at each token position in a sequence, where $\mathbf{x}_{t}$ represents the $t$th token in the output sequence. This allows them to generate a wide range of outputs, including natural language text, programming code, structured instructions, and more.
Some of the latest foundation models have evolved from traditional LLMs into what are now called large reasoning models, such as DeepSeek-R1 [73]. These models are typically fine-tuned on top of LLMs using reinforcement learning to promote step-by-step reasoning or tree-search-like problem-solving behaviors, often with a focus on enabling reflection and self-correction. DeepSeek-R1, for example, is built using group relative policy optimization (GRPO) training guided by feedback from DeepSeek-V3 [72]’s judgment feedback. Closely related are mathematical LLMs, such as rStar-Math [74], DeepSeek-Prover-V2 [75], MINIMO [76], and LLMOPT [77], which specialize in mathematical reasoning tasks. These models often adapt general reasoning methods to math problems, commonly incorporating inference-time scaling (akin to extending thinking time in humans) and leveraging tree search to increase logical depth.
In parallel, some models aim to fundamentally rethink the architecture and training strategies of LLMs, particularly through diffusion-based language models and world models. LLaDA [78] and MMaDA [79] use diffusion processes instead of autoregressive generation, enabling controllable output lengths, denoising-style masked generation, and multi-modal input-output support. World models, on the other hand, aim to learn latent representations of physical environments and predict object dynamics within them. Notably, the latest V-JEPA 2 [80] is pre-trained on large-scale unlabeled video data and then fine-tuned on a small set of robotic control videos, enabling V-JEPA 2 AC to support zero-shot robotic control.
These recent advancements in reasoning models, diffusion language models, and world models hold promising implications for network optimization tasks. However, how to effectively adapt and tailor them to networking scenarios remains an open challenge. In this stage, LPTMs have been increasingly applied across diverse scenarios, including mathematical modeling [81], analysis of optimization problems [54], and even approximate problem-solving [82].
@@ II-B3 Others
GenAI is not only applied to the above two types of generative models, but also has applications in other fields. Beyond images and text, GenAI has continued to advance across various multi-modal domains, including image-text contrastive learning (CLIP [83]), visual segmentation (SAM [84]), tabular data (CARTE [85]), graph-structured data (GraphEdit [86]), and multi-modal composition (CoDi [87]). These methods have also provided many new ideas for network optimization, especially the issue of multi-modal composition in the network, which has received much attention.
@@ II-C Neural Combinatorial Optimization
Combinatorial optimization refers to the class of problems in which the goal is to find the optimal configuration or sequence from a discrete, often combinatorially large, solution space, under given constraints. Classical examples include the traveling salesman problem (TSP), maximum independent set (MIS), and vehicle routing, all of which are NP-hard and difficult to scale with traditional solvers. Combinatorial optimization can be viewed as a class of problems encompassed by network optimization.
Neural combinatorial optimization aims to tackle these problems using neural networks to learn heuristics or direct solution mappings. GDMs have recently emerged as a powerful and popular approach for solving challenging combinatorial optimization problems [88, 89, 90, 91, 92]. These studies primarily focus on large-scale instances of classical combinatorial optimization problems, often involving hundreds of thousands of nodes, and aim to develop pre-trained models on canonical problem types that can generalize effectively to problem variants. Unlike reinforcement learning-based solvers [93], GDMs can directly model the solution distribution and generate near-optimal candidates in a single inference, making them well-suited for one-shot optimization scenarios common in network settings. This growing body of work has offered valuable inspiration for how GenAI, especially GDMs, can be applied to generate high-quality solutions in network optimization contexts.
Figure 4: Comparison between GDMs and LLMs in terms of generation mechanism. The sources of randomness in GDMs and LLMs differ: in GDMs, randomness arises from the initial noise sample and the stochastic nature of the denoising trajectory, whereas in LLMs, it stems from sampling predicted tokens. Moreover, LLMs are autoregressive, as each predicted token is fed back into the model for subsequent predictions, while most GDMs are non-autoregressive.
@@ II-D Deep Reinforcement Learning
DRL combines deep learning’s perception capabilities with reinforcement learning’s decision-making capabilities. It involves learning optimal policies through interactions between an agent and its environment to maximize cumulative rewards. By utilizing deep neural networks as function approximators, DRL is capable of addressing complex decision-making problems [94, 95].
GDMs are one of the earliest GenAI techniques applied in DRL, showcasing significant potential in robot control applications [96, 97, 98, 99, 100, 101, 102, 103, 104, 105]. Its unique capabilities in modeling complex data distributions have since enabled gradual expansion to broader DRL tasks. Recent advancements highlight GDM’s effectiveness in trajectory generation, providing smooth, diverse motion paths for agents [98]. In multi-task learning scenarios, GDMs facilitate knowledge transfer across different tasks, improving overall learning efficiency [105]. Moreover, through data augmentation, GDMs enhance the quality and diversity of training datasets, which is crucial for improving the generalization ability of DRL agents [99]. The application of GDMs in DRL is being widely migrated to the MDP framework of network optimization modeling [40]. Ongoing innovations in the AI community are driving performance improvements in MDP-based network optimization.
@@ III GDMs in Network Optimization
Most applications of GDMs in network optimization are inspired by advances in the AI community. These applications primarily fall into two categories: using GDMs as solution generators for combinatorial optimization problems [88, 89, 90, 91, 92], and using GDMs as policy networks in DRL for robotic control [96, 97, 98, 99, 100, 101, 102, 103, 104, 105], as shown in Table VI.
When used as solution generators for combinatorial optimization, GDMs are typically discrete diffusion models targeting large-scale problems (often involving over 100,000 nodes). In the AI domain, GDMs have achieved near-optimal performance on classical combinatorial optimization tasks such as TSP and MIS. Recent efforts have focused on accelerating solution generation [89] and enabling cross-problem generalization by leveraging shared knowledge across problem instances [92]. In the DRL setting, GDMs are commonly applied to generate action-state trajectories for robotic control. In robotic control, GDMs are mostly used in offline reinforcement learning, where they learn to predict the target trajectory from the history of trajectories observed during interaction with the environment. Most existing approaches execute only the first action of the generated trajectory, leading to inefficient exploration and significant time overhead. Current research is therefore exploring ways to refine generated trajectories through planning to improve action-generation efficiency [101, 104]. General-purpose DRL using GDMs has also been explored as a promising research direction [99, 100, 105].
@@ III-A GDMs as Solution Generators
Network optimization problems are inherently more diverse and complex than classical combinatorial optimization. They encompass not only discrete-variable problems but also non-convex optimization, mixed-integer programming, and related areas. Compared to traditional combinatorial tasks, network optimization problems often involve more intricate objective functions and constraints, as they must account for various factors such as connectivity, latency, and total available resources.
@@
GDMOPT [40] is one of the earliest works to apply GDMs as solution generators in the network domain, conducting a case study on a classical convex optimization problem. DiffSG [106] demonstrates the effectiveness and cross-domain generalization of GDMs on a variety of more complex network optimization tasks. [107] further provides a theoretical explanation of why GenAI can outperform discriminative models in solution generation, offering insights into GDMs’ underlying mechanisms and analyzing the impact of key hyperparameters in network optimization settings, with DiffSG as a baseline. DFSS [108] aims to address the security of channel state information (CSI)-based sensing in integrated sensing and communication (ISAC) scenarios. It tackles unauthorized sensing or spoofing attacks by adversaries trying to exploit CSI signals. DFSS first uses a discrete graph diffusion model to generate connection topologies for ISAC device deployment, minimizing operational costs while preserving sensing performance. Then, it applies a continuous diffusion model to generate safeguarding CSI values, embedding them into pilot signals to significantly reduce the sensing accuracy of unauthorized devices.
@@
GDSG [109] considers the common challenge in network optimization where optimal solutions are difficult to obtain. It extends GDM training to support robust learning from sub-optimal solution samples only. Trained solely on efficiently constructed sub-optimal datasets, GDSG successfully optimizes multi-server multi-user computation offloading problems in multi-access edge computing (MEC) scenarios, involving discrete diffusion for offloading decisions and continuous diffusion for resource allocation. DEITSP [110] focuses on reducing the computational complexity of autoregressive GDM solvers. DEITSP employs a one-step diffusion mechanism that enables inference from any timestep toward the target solution. It also introduces a novel modification to the DDIM [60] process: each solution candidate undergoes a one-step prediction, followed by renoising and reintegration into the denoising process, to enhance exploration. Additionally, it streamlines the GAT backbone to reduce inference time while maintaining solution quality, achieving faster results compared to DIFUSCO [88] and T2T [90]. GDPlan [111] addresses network planning by incorporating an external energy score-based model to guide graph diffusion for Internet Protocol (IP) link capacity configuration. High-quality solutions are further refined through external solvers in GDPlan.
@@
TABLE VI: Representative works on GDMs as solution generators and DRL policies for the AI community.
Type Algorithm Neural Backbone Contributions
@@
Energy-Guided Sampling [92] GCN Proposed a general energy-guided sampling framework that enables a pre-trained GDM for the TSP to generate solutions across TSP variants.
@@
IDQL [97] MLP Demonstrated that the commonly used simple Gaussian policy struggles to perform well in learning the Q-function solely from offline datasets under implicit Q-learning (IQL) settings. At the same time, GDMs can effectively take on this role with their powerful multimodal distribution-learning capability.
@@
MTDIFF [99] Transformer Leveraged GDMs for both trajectory planning and data synthesis in multi-task DRL.
@@
EDP [101] MLP Designed an action approximation technique that enables direct reconstruction of actions during training without executing the full sampling chain, combined with the DPM-Solver, significantly improving training efficiency.
SRDP [102] MLP Considered out-of-distribution (OOD) generalization issues in offline reinforcement learning caused by a mismatch between real-world states and the training distribution by using GDMs to learn generalizable state reconstruction features.
@@ III-B GDMs as DRL Policy Networks
Many network optimization problems can be modeled as MDP problems, but they differ from traditional robotic control tasks in two significant ways. First, the reward functions (or objective functions) and constraints in network optimization are typically far more complex than those in robotic control. While robot control tasks often have clear terminal states, making it easier to learn critic modules, network optimization problems often involve non-differentiable objectives and constraints, which cannot be directly optimized and must be approximated via external relaxation. These surrogate modules are sensitive to the objective formulation and hyperparameter settings. Second, in terms of real-time performance, dynamic network optimization typically requires decision-making within milliseconds to avoid system failures. In contrast, the actions generated by GDMs in robotic control are typically handled at higher-level planning layers, with real-time control handled at lower levels, thereby relaxing real-time constraints on GDMs.
@@
DLRS [121] addresses the use of vector databases for request caching in edge-deployed LLMs. Observing that existing methods overlook previous search results, [121] proposes a new MDP-based quality of service (QoS) optimization framework called DLRS, improving both similar and novel request response frequencies in a cloud-edge cooperative setting. In [122], a GDM serves as a contract generator in a downlink semantic communication system, jointly optimizing rate-splitting multiple access (RSMA)-based resource allocation and contract design, achieving higher rewards than traditional DRL and supporting downstream semantic segmentation tasks. AC-GDM [123] targets secure beamforming in a multi-user IoT system with imperfect channels, using GDM-based DRL to jointly optimize precoding at the base station and phase shifts at an intelligent reflecting surface (IRS), outperforming traditional methods in simulations. GAI-Asynchronous [124] considers a low-altitude multi-task over-the-air edge inference system, modeling the joint optimization of task batching and beamforming, and uses GDMs as DRL policies to perform batch-wise beamforming decisions.
In summary, GDMs demonstrate superior performance as DRL policy networks in MDP-formulated network-optimization problems, thanks to their powerful ability to model complex, multimodal action distributions. However, as in the broader AI community, the high latency of GDMs due to their iterative denoising process remains a limitation. Furthermore, the complexity of network optimization objectives and constraints presents substantial challenges for developing general-purpose GDM-based DRL approaches and achieving cross-problem generalization in the network domain.